3 4 Y 3 2 1 Consider thee equation Y 5 |0- Find the solution with initial condition Y (0)
Answer #1

9 26ACSt-Ret 2A-28 J-22 + -_ 2 Apr T 0G 49 C4-31 2 4-37 (T-50-6)2 1-31+ 3 0-3)0D-2)-123 3)t +3 IY D-T=h-1R 5- -3 t(D-2)2 t- T[48301부- 상e. 하용 G=CO)PR 2